# The Internet of Us

### v2

@@ 5 Supporting and Incentivizing the Self-Determination of Communities
Awareness of descriptive diversity and the dimensions along which it is measured supports the self-determination of communities by providing the primitive concepts against which a community can differentiate itself from other communities. A community could, for instance, self-determine as including only residents of Spain who are wine aficionados, effectively utilizing the “country of residence” and the “hobbies and leisure” features of its members as the properties that make it distinct from other communities.
@@
The same primitive concepts support a second path to community self-determination, by providing the basis over which a community can define the norms that apply during the interaction between its members. Sociologically speaking, norms capture the rules of engagement, or the social rules of behavior of a community and its members, and are the direct application of the social values of that community. They capture community-wide provisions (e.g., do not ask the same person for support more than three times within any single week) or individual provisions (e.g., do not share a request for support with members blocked by an individual). Norms are also meant to interpret the search directives provided by members of a community. The same cue (e.g., seeking support from people nearby) can be interpreted differently if seeking support within one’s “neighborhood watch” community, or within one’s “alma mater graduates” community. Technologically speaking, norms define the protocol of exploring the search space when seeking to solve the search problem of identifying appropriate respondents to a sequence of queries.
@@
Norms come in the form of natural language expressions in human societies, and, traditionally, in the form of logic-based expressions in formal machine-readable representations (e.g., constraint-based [28], time-based [49], or event-based [50]). In our work, we have adopted a declarative representation of norms through conditional logic-based rules [17]. The rule premises refer to profile features (such as gender, competencies, etc.) or other state properties of the platform (such as the number of active community members, the total number of queries, the level of user satisfaction, etc.) that can encode information about a community, a user, or even a task or interaction. The rule conclusions refer to the possible actions that can be performed, which include message-passing actions (such as sending a message to another user, or sending a message to an app’s user-facing interface to enable a given button), and actions that change profile features or state properties (such as the privacy setting of one’s location, or the rating for a given task). A norm engine is then designed to mediate interactions by checking their compliance with the relevant norms in place, be they community-wide norms or individual norms.
@@
Of course, having community members aligned in their understanding of norms is crucial, but difficult in open systems. For instance, if a norm prohibits hate speech, not all members might have the same understanding of what constitutes hate speech. We have developed learning mechanisms that can learn the meaning of a norm from a community’s interactions, explain this meaning to community members, and adapt over time based on members’ feedback [51]. Furthermore, working with norms allows us to formally analyze the outcomes of interactions [52]. For instance, we can analyze whether certain norms promote/demote certain human values [53], as well as synthesize norms to ensure the maximum promotion of given human values [54].
@@
How can we promote the growth of active, inclusive, and diverse communities, taking also into account that the platform allows for interactions with unknown unknowns? In the spirit of community self-determination, the key consideration is that communities may not wish for external guidelines on inclusiveness and active participation to be applied in a paternalizing and all-knowing manner. Consequently, instead of imposing such guidelines, our platform seeks to incentivize their adoption, taking into account the diverse characteristics of different participants and communities, and avoiding a “one-size-fits-all” approach common in current systems. Past works have recognized user diversity and harnessed it for generating incentives using badges [25], text messages [55], and recommended items [56]. Recently, we proposed a diversity-aware multi-armed bandit approach [27, 26], which reasons about users’ diversity characteristics when making action recommendations. Such reasoning was shown to lead to improved outcomes compared to a “static” approach, which ignored participants’ rich and changing profiles (cf. Section 3).
@@
The platform’s incentivization mechanism directs participants to engage with others by seeking support, responding to support queries, and rating the responses they receive, towards becoming more active members of their community. This same mechanism could be used to incentivize participants not to apply search filters unnecessarily, and to actively request that the recipients of their query be diversified (cf. Section 6).
@@ 6 Modeling the Personal and the Social Context of Individual Participants
One of the aspects that makes individuals diverse, and that affects their social interactions through the application of norms, is their personal context [57, 58]. Due to its inherently dynamic nature, the personal context of any given platform participant cannot be populated in their profile when they first join the platform, but needs to be learned. The worldwide deployment of diversity pilots over our platform presented a unique opportunity to investigate how personal contexts can be learned across diverse communities. Short surveys about the participants’ current place, activities, social situation, and mood were collected in parallel with data from smartphone sensors and mobile app logs. The datasets were used to investigate how diversity attributes, such as the country of residence of participants, can be adopted to build ML-based inference models from mobile data. Such an investigation allows for the systematic comparison of country-specific and country-agnostic models for tasks like activity recognition [29] or mood recognition [47], and facilitates the empirical evaluation of whether model transfer is possible across communities.
@@
The results indicate that considering country-level diversity in model learning is crucial for recognition performance, and transferring models between countries is not a simple task. In this sense, the results contribute towards highlighting the fundamental need for locally-valid data that represents the reality of the world beyond economically advantaged regions [59], and for documenting the resulting models with respect to these issues [60]. More broadly, our empirical results emphasize the need for machine learning models to be aware of the diversity that exists between communities, and by extension, the need for this diversity to be reflected in the learning data that is available (in the participant profiles, in the case of our work).
@@
Beyond their personal context, a profile also includes an individual’s social context, which in its simplest form comprises a list of learnable social tie strengths [30] that capture the frequency and manner in which the individual has interacted with other platform participants. A more advanced form of social context comes through a set of norms that specify an individual’s policy on whether / how to interact with other participants. This policy can be thought to be more persistent than an individual’s social tie strengths, but can still be revised over time. Our work has focused on developing a formal declarative policy-representation language [61] that acknowledges the key role of argumentation as a calculus for human-centric AI [32]. A policy elicitation paradigm accompanies the language [33, 34] that is cognitively light and computationally efficient [31]. To support the utilization of this formal policy language, we have investigated how policies can be elicited through natural language cues [62], and how they can be integrated with machine learning models [63], towards making them explainable [64] and contestable [35].
@@
The use of individual norms to condition a user’s interactions on their personal and social context, and the dependence of the user’s context on their interactions, may lead to a vicious spiral of diversity-reducing interactions. To compensate for this possibility, we adopt a second view of diversity as a prescriptive value to be promoted, and the platform diversifies the ranking of participants that receive any given query. One might think that diversification should be applied maximally on all visible characteristics, but not on any of the invisible ones, which relate to the abilities of participants to respond to queries and should, therefore, be allowed to be constrained. This view is understandable, given that human societies have practically equated the notion of diversity with that of inclusiveness based on visible characteristics. But this, by itself, does not make this a valid view.
@@
In certain scenarios, one may wish to select participants based on their visible characteristics, to target, for instance, a particular minority demographic group. In other scenarios, one may want to diversify based on the participants’ invisible characteristics, not necessarily as a way to avoid discrimination, but as a way to get a more varied perspective (or serendipity, cf. Section 4). In the spirit of self-determination, our diversification process allows each community to choose the profile characteristics (one or many, visible or invisible) against which to diversify, and the extent to which the diversity requirement will override a user’s own preferences [38]. Empirical studies on how humans perceive diversity in participant rankings have been carried out to validate the process.
@@ 7 Diversity as an Instrumental Value that Supports Fundamental Values
Any attempt to operationalize and mediate diversity in social interactions must consider the possible ethical ramifications. Technological mediation is not value-neutral, and the development of technology is never done without prioritizing different values and advancing a certain “agenda” [65, 66, 42]. Being mindful of the inevitably value-ridden nature of the technology development process, the IoU platform seeks to promote the participants’ exposure to diversity to the extent that it does not pose a threat to their safety. We, thus, have to carefully weigh the values of inclusion and protection to enable a meaningful exposure to diversity [67].
Our rationale follows the assumption that diversity is not an end, but a means; it has instrumental rather than intrinsic value [39]. In the same way that diversity may promote the value of inclusiveness, it may also, if abused, demote the value of protection. This relates to the classic discussion in content moderation: the need to balance the freedom of speech (as promoted in liberal democracies) and the prevention of hate speech and violence [68, 69]. On the one hand, public discourse should include as many perspectives as possible. On the other hand, public discourse should be free from abusive communication (such as racist, sexist, and antisemitic attacks) and protect minority positions.
Values like inclusiveness and protection are of a more fundamental nature vis à vis diversity [40], and they ought to be promoted for their own sake, even at the expense of diversity. Think of a request about a topic that is associated with stigma (e.g., mental health problems), where an individual may risk social exclusion just for seeking support [41]. Somewhat counter-intuitively, restricting the diversity of the respondents ends up enhancing the individual’s access to a diverse section of the social search space, by protecting them from hurtful and derogatory comments, and by creating a safe space within which to seek support. Accordingly, our platform implements algorithms and offers interface functionalities to interact anonymously and to restrict the respondent pool. At the same time, transparency is essential to enable individual control over personal data and to trust in the potential of a diversity-aware platform like WeNet [42].
@@ 8 Conclusion
We have sought to make progress on an ambitious goal: to design the IoU and an AI-mediated and diversity-aware platform for social interactions enabling it. In attempting to fulfill that goal, we have explored various facets of diversity while, notably, staying grounded in empirical work across a diverse set of participating communities. Among our contributions is the gathering and eventual sharing of social interaction data at a scale that goes beyond past attempts. This is only a first step, and there is still a long way to go. But we believe that the direction of rethinking social interactions under a diversity paradigm is promising, and it does raise several relevant research questions. Ultimately, the main take-home message is the importance of a holistic, empirically-grounded, and multidisciplinary perspective in pushing towards a future of diversity-aware hybrid human-AI societies.
It goes without saying that realizing the IoU vision raises several key issues that this paper has not even mentioned. A key issue is how to make people searchable and findable while maintaining their privacy. This is crucial because people must disclose specific properties, such as their location and skills, which may ultimately make them recognizable. A second issue is security - how to manage malicious behavior. A third is how to keep people motivated in providing help to others. The work done so far is only a very first preliminary step. A fourth related issue is how to minimize disturbance. People would definitely lose motivation if they were asked for intervention too often. A fifth issue is data and profile quality. Data that changes frequently is costly to maintain. And so on. No IoU can be enabled without a solid solution to these issues. At the same time, this work provides a first insight into the issue of how to understand and manage diversity, an issue which we believe is foundational and that, we believe, had not been previously studied with the necessary depth.
@@ Acknowledgements
This research has received funding from the European Union’s Horizon 2020 FET Proactive project “WeNet — The Internet of Us”, under grant agreement no. 823783. We wish to thank all our collaborators for their fantastic support; a most likely incomplete list is: Marina Bidoglia, Andrea Bontempelli, Maria Chiara Campodonico, Carlo Caprini, Ronald Chenu Abente, Galileo Disperati, Paula Helm, Alethia Hume, Peter Kun, Meegahapaola Lakmal, Vasileios Theodoros Markos, Chaitanya Nutakki, Marcelo Dario Rodas Brites, and Donglei Song.

### v3

@@ 4 Supporting and Incentivizing the Self-Determination of Communities
Awareness of descriptive diversity and the dimensions along which it is measured, supports the self-determination of communities by providing the primitive concepts against which a community can differentiate itself from other communities. A community could, for instance, self-determine as including only residents of Spain who are wine aficionados, effectively utilizing the “country of residence” and the “hobbies and leasure” features of its members as the properties that make it distinct from other communities.
@@
The same primitive concepts support a second path to community self-determination, by providing the basis over which a community can define the norms that apply during the interaction between its members. Sociologically speaking, norms capture the rules of engagement, or the social rules of behavior of a community and its members, and are the direct application of the social values of that community. They capture community-wide provisions (e.g., do not ask the same person for support more than three times within any single week), or individual provisions (e.g., do not share a request for support with members blocked by an individual). Norms are also meant to interpret the search directives provided by members of a community. The same cue (e.g., seek support from people nearby) can be interpreted differently if seeking support within one’s “neighborhood watch” community, or within one’s “alma mater graduates” community. Technologically speaking, norms define the protocol of exploring the search space when seeking to solve the search problem of identifying appropriate respondents to a sequence of queries.
@@
Norms come in the form of natural language expressions in human societies, and, traditionally, in the form of logic-based expressions in formal machine-readable representations (e.g., constraint-based [Garcia-CaminoRSV09], time-based [AgotnesHRSW09], or event-based [KowalskiS05]). In our work, we have adopted a declarative representation of norms through conditional logic-based rules [OsmanCSSG21]. The rule premises refer to profile features (such as gender, competencies, etc.) or other state properties of the platform (such as the number of active community members, the total number of queries, the level of user satisfaction, etc.) that can encode information about a community, a user, or even a task or interaction. The rule conclusions refer to the possible actions that can be performed, which include message-passing actions (such as sending a message to another user, or sending a message to an app’s user-facing interface to enable a given button), and actions that change profile features or state properties (such as the privacy setting of one’s location, or the rating for a given task). A norm engine is then designed to mediate interactions by checking their compliance with the relevant norms in place, be they community-wide norms or individual norms.
@@
Of course, having community members aligned in their understanding of norms is crucial, but difficult in open systems. For instance, if hate speech is prohibited by a norm, not all members might have the same understanding of what constitutes hate speech. We have developed learning mechanisms that can learn the meaning of a norm from a community’s interactions, explain to community members that learned meaning, and adapt over time what is learned through members’ feedback [SantosOS22]. Furthermore, working with norms allows us to formally analyze the outcomes of interactions [MontesOS22]. For instance, we can analyze whether certain norms promote / demote certain human values [abs-2110-09240], as well as synthesize norms to ensure the maximum promotion of given human values [MontesS21].
@@
How can we promote the growth of active, inclusive, and diverse communities, taking also into account that the platform allows for interactions with unknown unknowns? In the spirit of community self-determination, the key consideration is that communities may not wish for external guidelines on inclusiveness and active participation to be applied in a paternalizing and all-knowing manner. Consequently, instead of imposing such guidelines, our platform seeks to incentivize their adoption, taking into account the diverse characteristics of different participants and communities, and avoiding a “one-size-fits-all” approach common in current systems. Past works have recognized user diversity and harnessed it for generating incentives using badges [yanovsky2019one], text messages [segal2018optimizing], and recommended items [ben2022generating]. Recently, we proposed a diversity-aware multi-armed bandit approach [botta2023sequencing, botta2023personalizing], which reasons about users’ diversity characteristics when making action recommendations. Such reasoning was shown to lead to improved outcomes compared to a “static” approach which ignored participants’ rich and changing profiles (cf. Section 2).
@@
The platform’s incentivization mechanism directs participants to engage with others by seeking support, responding to support queries, and rating the responses they receive, towards becoming more active members of their community. This same mechanism could be used to incentivize participants to not apply search filters unnecessarily, and to actively request for the recipients of their query to be diversified (cf. Section 5).
@@ 5 Modeling the Personal and the Social Context of Individual Participants
One of the aspects that makes individuals diverse, and that affects their social interactions through the application of norms, is their personal context [KD-2017-PERCOM, bontempelli2022lifelong]. Due to its inherently dynamic nature, the personal context of any given platform participant cannot be populated in their profile when they first join the platform, but needs to be learned. The world-wide deployment of diversity pilots over our platform presented a unique opportunity to investigate how personal contexts can be learned across diverse communities. Short surveys about the participants’ current place, activities, social situation, and mood were collected in parallel with data from smartphone sensors and mobile app logs. The datasets were used to investigate how diversity attributes like the country of residence of participants can be utilized to build ML-based inference models from mobile data. Such an investigation allows for the systematic comparison of country-specific and country-agnostic models for tasks like activity recognition [Bouton22] or mood recognition [10.1145/3569483], and facilitates the empirical evaluation of whether model transfer is possible across communities.
@@
The results show that taking country-level diversity into account as part of model learning is important from the perspective of recognition performance, and that transfer of models across countries is not a straightforward task. In this sense, the results contribute towards highlighting the fundamental need for locally-valid data that represents the reality of the world beyond economically advantaged regions [GaticaPerez19], and for documenting the resulting models with respect to these issues [10.1145/3458723]. More broadly, our empirical results emphasize the need for machine learning models to be aware of the diversity that exists between communities, and by extension, the need for this diversity to be reflected in the learning data that is available (in the participant profiles, in the case of our work).
@@
Beyond their personal context, a profile also includes an individual’s social context, which in its simplest form comprises a list of learnable social tie strengths [Perikos.2022] that capture the frequency and manner in which the individual has interacted with other platform participants. A more advanced form of social context comes through a set of norms that specify an individual’s policy on whether / how to interact with other participants. This policy can be thought to be more persistent than an individual’s social tie strengths, but can still be revised over time. Our work has focused on developing a formal declarative policy-representation language [Markos.2022.prudens] that acknowledges the key role of argumentation as a calculus for human-centric AI [Dietz.2022.calculus]. The language is accompanied by a policy elicitation paradigm [Michael.2019.coaching, Michael.2023.coachable] that is cognitively light and computationally efficient [Markos.2022.proxy]. To support the utilization of this formal policy language, we have investigated how policies can be elicited through natural language cues [Ioannou.2021.translation], and how they can be integrated with machine learning models [Tsamoura.2021], towards making them explainable [Michael.2023.fourthAI] and contestable [Tubella.2020].
@@
The use of individual norms to condition a user’s interactions on their personal and social context, and the dependence of the user’s context on their interactions may lead to a vicious spiral of diversity-reducing interactions. To compensate for this possibility, we adopt a second view of diversity as a prescriptive value to be promoted, and the platform diversifies the ranking of participants that receive any given query. One might think that diversification should be applied maximally on all visible characteristics, but not on any of the invisible ones, which relate to the abilities of participants to respond to queries and should, therefore, be allowed to be constrained. This view is understandable, given that human societies have practically equated the notion of diversity with that of inclusiveness based on visible characteristics. But this, by itself, does not make this a valid view.
@@
In certain scenarios, one may wish to select participants based on their visible characteristics, so as to target, for instance, a certain minority demographic group. In other scenarios, one may wish to diversify based on the participants’ invisible characteristics, not necessarily as a way to avoid discrimination, but as a way to get a more varied perspective (or serendipity, cf. Section 3). In the spirit of self-determination, our diversification process allows each community to choose the profile characteristics (one or many, visible or invisible) against which to diversify, and the extent to which the diversity requirement will override a user’s own preferences [Markos.2022.diversity]. Empirical studies on how humans perceive diversity in participant rankings have been carried out to validate the process.
@@ 6 Diversity as an Instrumental Value that Supports Fundamental Values
Any attempt to operationalize and mediate diversity in social interactions must consider the possible ethical ramifications. Technological mediation is not value-neutral, and the development of technology is never done without prioritizing different values and advancing a certain “agenda” [Friedman.1996, Brey.2010, schelenz2024transparency]. Being mindful of the inevitably value-ridden nature of the technology development process, our platform — born out of a European, democratic, and liberal context — seeks to promote the participants’ exposure to diversity to the extent that it does not pose a threat to their safety. We, thus, have to carefully weigh the values of inclusion and protection to enable a meaningful exposure to diversity.
Our rationale follows the assumption that diversity is not an end, but a means; it has instrumental rather than intrinsic value [Zimmerman.2019]. In the same way that diversity may promote the value of inclusiveness, it may also, if abused, demote the value of protection. This relates to the classic discussion in content moderation: the need to balance the freedom of speech (as promoted in liberal democracies) and the prevention of hate speech and violence [Gillespie.2018, Brown.26February2021]. On the one hand, public discourse should include as many perspectives as possible. On the other hand, public discourse should be free from abusive communication (such as racist, sexist, and antisemitic attacks) and protect minority positions.
Values like inclusiveness and protection are of a more fundamental nature vis à vis diversity [Helm.23.09.2021, Helm.2022], and they ought to be promoted for their own sake, even at the expense of diversity. Think of a request about a topic that is associated with stigma (e.g., mental health problems), where an individual may risk social exclusion just for seeking support [Goffman.1990]. Somewhat counter-intuitively, restricting the diversity of the respondents ends up enhancing the individual’s access to a diverse section of the social search space, by protecting them from hurtful and derogatory comments, and by creating a safe space within which to seek support. Accordingly, our platform implements algorithms and offers interface functionalities to interact anonymously and to restrict the respondent pool.
@@ 7 Conclusion
We have sought to make progress on an ambitious goal: to design an AI-mediated and diversity-aware platform for social interactions. In attempting to fulfill that goal, we have explored various facets of diversity, while, importantly, staying grounded in empirical work across a diverse set of participating communities. Among our contributions is the gathering, and the eventual sharing, of social interaction data at a scale that goes beyond past attempts. This is only a first step, and there is still a long way to go. But we believe that the direction of rethinking social interactions under a diversity paradigm is promising, and it does raise several relevant research questions. Ultimately, the main take-home message is the importance of a holistic, empirically-grounded, and multidisciplinary perspective in pushing towards a future of diversity-aware hybrid human-AI societies.
@@ Acknowledgements
This research has received funding from the European Union’s Horizon 2020 FET Proactive project “WeNet — The Internet of Us”, under grant agreement no. 823783. We wish to thank all our collaborators for their amazing support; a, most likely, incomplete list is: Marina Bidoglia, Andrea Bontempelli, Maria Chiara Campodonico, Carlo Caprini, Ronald Chenu Abente, Galileo Disperati, Paula Helm, Alethia Hume, Peter Kun, Meegahapaola Lakmal, Vasileios Theodoros Markos, Marcelo Dario Rodas Brites, and Donglei Song.